3. An ideal shock absorption system would use a critically damped oscillator to absorb shock loads. The location of the absorbing piston (π‘₯) is described by π‘₯ = πœπ‘’βˆ’π›Ύπ‘‘ where:

- 𝜏 is the linear damping coefficient

- 𝛾 is the exponential damping constant

- 𝑑 is the time (𝑠)

- π‘₯ is the displacement of piston (π‘š)

The tasks are to:

a) Draw a graph of displacement against time for 𝜏 = 12 and 𝛾 = 2, between 𝑑 = 0𝑠 and 𝑑 = 10𝑠.

b) Calculate the gradient at 𝑑 = 2𝑠 and 𝑑 = 4𝑠.

Β QD099_September_2017

Page 7 of 10


Β Β Β c) Differentiate the function of π‘₯ and calculate the value of 𝑑π‘₯ at 𝑑 = 2𝑠 and 𝑑 = 4𝑠. 𝑑𝑑

d) Compare your answers for part b and part c. (M1)

e) Calculate the derivative for the velocity function(𝑑2π‘₯).


1
Expert's answer
2022-03-10T13:21:16-0500

Solution


Given that


"x = \tau e - \gamma t\"


Now for "\tau = 12,\,\,\,\gamma = 2\" , we have


"x = 12e - 2t\"


The plot between "t = 0" and "t = 10" isΒ 




Solution (b)


Since the plot is a straight line, therefore, the gradient is constant which is "m=-2"


Henec gradient when "t=2" s is "m=-2"


And gradient when "t=4" s is "m=-2"


Solution (c)


"x = 12e - 2t\"


"\frac{{dx}}{{dt}} = 0 - 2(1) = - 2\"


Hence "t=2" s , we have gradient "\frac{{dx}}{{dt}} = - 2\"


And

Hence "t=4" s , we have gradient "\frac{{dx}}{{dt}} = - 2\"



Solution (d)


From (b) and (c), we see that the gradients when "t=2" s and when "t=4" s is "m=-2", fix that is the same.



Solution (e)


"x = 12e - 2t\"


velocity is "v=\frac{{dx}}{{dt}} = 0 - 2(1) = - 2\"